OpenStudy (anonymous):

indicate wether the given series is converge or diverge if it converges find the sum from zero to infinity (1)/(sqrt 7)^k

OpenStudy (anonymous):

Sorry lol. Nt very gud in this

OpenStudy (anonymous):

in converges for sure

OpenStudy (anonymous):

\[\sum_0^{\infty}\frac{1}{\sqrt{7}^k}\] converges because \[|\frac{1}{\sqrt{7}}|<1\]

OpenStudy (anonymous):

use \[\sum_{k=0}^{\infty}x^k =\frac{1}{1-x}\] \[\sum_{k=0}^{\infty}\frac{1}{\sqrt{7}^k}=\frac{1}{1-\frac{1}{\sqrt{7}}}\]